what is the mass percent of 20.0grams Mg(NO3)2 dissolved in 500.0 grams of water?
Answer #1

% mass = (mass of solute / mass of solution) x 100

mass of solute = 20.0 g

total mass = 500 + 20 = 520 g

% mass = (20 / 520) x 100

% mass = 3.85 %